# Complex Domain Approach for Reversible Data Hiding and Homomorphic Encryption: General Framework and Application to Dispersed Data

### v1

@@ III-C Data Aggregation
$\sigma^{\prime}_{k}=\lambda_{k}\sum_{j=1}^{N}\delta_{j,k}=\lambda_{k}\sum_{j=1}^{N}(d_{j,k}+iw_{k}).$
By separating the real and imaginary components of the sum, and noting that the watermark $w_{k}$ is the same for all devices in a given round, the expression for $\sigma^{\prime}_{k}$ can be rearranged as:
@@
The extraction of both the aggregated data $S_{k}$ and the watermark $w_{k}$ from $\sigma^{\prime}_{k}$ is straightforward. We can first multiply $\sigma^{\prime}_{k}$ by the inverse of $\lambda_{k}$ in the complex domain, $\lambda_{k}^{-1}$. Then, the real part of the result directly yields the aggregated original data ($S_{k}$), and the imaginary part, divided by the number of devices ($N$), gives the watermark ($w_{k}$). This extraction process can be formally expressed as:
@@
A numerical example with small numbers is provided in the Appendices (supplemental material) to illustrate how the embedding and recovery methods work, both for single data and for data aggregration.
@@
IV Homomorphic Encryption on the Complex Domain
@@
For a better understanding of the adaptation of this cryptosystem to the complex domain, a brief introduction to modular arithmetic over Gaussian integers is provided in the Appendices (supplemental material). A more detailed reference on Gaussian integers can be found in [7].
IV-A ElGamal Encryption over Gaussian Integers
@@
We define the “Gaussian” version of ElGamal encryption as follows. Let $p$ be a Gaussian prime (preferably a regular prime integer of the form 4n14n-1, as discussed in Appendix A). Let $\gamma$ be a generator of the cyclic multiplicative group of Gaussian integers modulo $p$, denoted as ${\mathbb{Z}[i]}^{*}_{p}$. The ElGamal encryption scheme over Gaussian integers consists of the same three fundamental algorithms:
@@
As a key property, the order of this group, $\mathcal{N}=p^{2}-1$, is significantly larger than p1p-1 for the same numerical prime $p$ used in the regular integer ElGamal. This relationship, $\mathcal{N}=(p-1)(p+1)$, is particularly advantageous for security. If a prime $p$ is chosen such that p1p-1 possesses a large prime factor (a standard requirement for robust security in integer-based ElGamal against algorithms like Pohlig-Hellman [18]), then this same large prime factor will inherently be a factor of $\mathcal{N}$. Therefore, the Gaussian integer ElGamal variant naturally inherits at least the same level of security (in terms of resistance to Pohlig-Hellman attacks) as its regular integer counterpart for a given $p$. Furthermore, the existence of the additional factor $(p+1)$ in $\mathcal{N}$ can potentially introduce even larger prime factors, further enhancing the security of the scheme. Thus, careful selection of the Gaussian prime $p$ to ensure that the largest prime factor of $\mathcal{N}$ is sufficiently large guides robust parameter choices.
@@ IV-A1 Numerical example
For the parameters, the prime modulus $p=23$ and the generator $\gamma=1+2i$ are taken, and the private key is chosen as $a=7$. Now, $\mathcal{K}$ is computed as $\mathcal{K}=\gamma^{a}\,(\operatorname{\mathrm{mod}}p)=(1+2i)^{7}\,(\operatorname{\mathrm{mod}}23)=6+2i$, and the public key is thus formed as the triple $\mathbf{K}=(p,\gamma,\mathcal{K})=(23,1+2i,6+2i)$.
To encrypt a first message $\mu_{1}=5+4i$, an ephemeral key $b_{1}=5$ is chosen. The ciphertext components are computed as $\psi_{1}=\gamma^{b_{1}}\,(\operatorname{\mathrm{mod}}p)=(1+2i)^{5}\,(\operatorname{\mathrm{mod}}23)=18+8i$, and $\psi_{2}=\mu_{1}\cdot\mathcal{K}^{b_{1}}\,(\operatorname{\mathrm{mod}}p)=(5+4i)\cdot(6+2i)^{5}\,(\operatorname{\mathrm{mod}}23)=21+11i$. The resulting ciphertext is $(\psi_{1},\psi_{2})=(18+8i,21+11i)$. Decryption of this ciphertext using the private key $a=7$ successfully yields $\mu_{1}=5+4i\,(\operatorname{\mathrm{mod}}23)$.
For a second message $\mu_{2}=3+2i$, an ephemeral key $b_{2}=7$ is chosen. The ciphertext components are computed as $\psi^{\prime}_{1}=\gamma^{b_{2}}\,(\operatorname{\mathrm{mod}}p)=(1+2i)^{7}\,(\operatorname{\mathrm{mod}}23)=6+2i$, and $\psi^{\prime}_{2}=\mu_{2}\cdot\mathcal{K}^{b_{2}}\,(\operatorname{\mathrm{mod}}p)=(3+2i)\cdot(6+2i)^{7}\,(\operatorname{\mathrm{mod}}23)=21+16i$. The resulting ciphertext is $(\psi^{\prime}_{1},\psi^{\prime}_{2})=(6+2i,21+16i)$. Decryption of this ciphertext using the private key $a=7$ successfully yields $\mu_{2}=3+2i\,(\operatorname{\mathrm{mod}}23)$.
@@
$\psi^{\prime\prime}_{1}$ $=\psi_{1}\cdot\psi^{\prime}_{1}=(18+8i)\cdot(6+2i)=0+15i\,(\operatorname{\mathrm{mod}}23),$
$\psi^{\prime\prime}_{2}$ $=\psi_{2}\cdot\psi^{\prime}_{2}=(21+11i)\cdot(21+16i)=12+15i\,(\operatorname{\mathrm{mod}}23).$
@@
Decryption of this combined ciphertext $(\psi^{\prime\prime}_{1},\psi^{\prime\prime}_{2})=(0+15i,12+15i)$ using the private key $a=7$ yields $\mu^{\prime\prime}=7+22i\,(\operatorname{\mathrm{mod}}23)$. This precisely matches the direct multiplication of the original plaintexts: $\mu_{1}\cdot\mu_{2}=(5+4i)\cdot(3+2i)\,(\operatorname{\mathrm{mod}}23)=(15-8)+(10+12)i\,(\operatorname{\mathrm{mod}}23)=7+22i\,(\operatorname{\mathrm{mod}}23)$. This confirms the correct operation and homomorphic property of ElGamal encryption over Gaussian integers.
IV-B Component-wise Paillier Encryption of Gaussian Integers
While the previously discussed ElGamal scheme was extended to operate natively over Gaussian integers, the additive homomorphic property of Paillier encryption allows for a straightforward application to complex numbers (specifically, Gaussian integers) without requiring a specialized “complex” version of the cryptosystem. This is achieved by encrypting the real and imaginary parts of a Gaussian integer independently.
@@ IV-B Component-wise Paillier Encryption of Gaussian Integers
$\mathrm{Dec}_{\mathsf{K_{priv}}}(c_{\mathrm{sum},R})$ $=m_{1R}+m_{2R}\,(\operatorname{\mathrm{mod}}n),$
$\mathrm{Dec}_{\mathsf{K_{priv}}}(c_{\mathrm{sum},I})$ $=m_{1I}+m_{2I}\,(\operatorname{\mathrm{mod}}n).$
@@
V Protocols for Joint Watermarking and Encryption on the Complex Domain
@@
The first protocol, termed the H[i][i]dden-EG protocol, uses Gaussian ElGamal encryption to provide joint watermarking and encryption for individual data points. The second one, denoted as the H[i][i]dden-AggP protocol, uses Paillier encryption’s additive homomorphism to enable joint watermarking, encryption, and privacy-preserving aggregation of data from multiple sources. Each protocol is presented with its initialization and data exchange phases, and also their respective complexities.
V-A H[i][i]dden-EG: Joint Watermarking and Encryption for Individual Data
@@
This section describes a protocol for secure and reversible information hiding between a sensor (S) and a data collector (DC). It applies the proposed complex domain framework and Gaussian ElGamal encryption to ensure sensor data ($d_{k}$) is transmitted confidentially while simultaneously embedding a time-dependent watermark ($w_{k}$) for verification and integrity.
The protocol proceeds in two distinct phases: Initialization and watermarking and data exchange.
@@
The initialization phase establishes the necessary cryptographic keys and shared secrets between the DC and S:
@@
    Watermark Agreement: DC and S agree on a set of shared watermarking keys or a mechanism to generate sequential, time-dependent watermarks: $w_{1},w_{2},\ldots,w_{M}$ corresponding to times $t_{1},t_{2},\ldots,t_{M}$. Each $w_{j}$ is a non-negative integer represented with $B$ bits, taking values in the range $0,\ldots,2^{B}-1$.
@@
    S then sends the resulting combined ciphertext $(\psi^{\prime\prime}_{k,1},\psi^{\prime\prime}_{k,2})$ to DC. This step simultaneously performs both embedding and encryption, with the additional benefit of homomorphic multiplication.
@@
    DC’s Decryption and Verification: Upon receiving the combined ciphertext $(\psi^{\prime\prime}_{k,1},\psi^{\prime\prime}_{k,2})$, DC decrypts it using its private key $a$. Due to the homomorphic property, the decrypted value $\mu_{k}=\mathrm{Dec}_{a}(\psi^{\prime\prime}_{k,1},\psi^{\prime\prime}_{k,2})$ will correspond to the product of the two original plaintexts: $\mu_{k}=\lambda_{k}\cdot\delta_{k}\,(\operatorname{\mathrm{mod}}p)$. Given the careful selection of $p$ to be sufficiently large, this modular result $\mu_{k}$ is guaranteed to be equivalent to the exact, non-modular product $\lambda_{k}\cdot\delta_{k}$ as Gaussian integers. To isolate the watermarked data, DC then computes $\delta_{k}=\mu_{k}/\lambda_{k}$ using standard complex number division. Finally, DC extracts the original sensor data $d_{k}$ from the real part of $\delta_{k}$ and the embedded watermark $w_{k}$ from its imaginary part. The data $d_{k}$ is accepted only if the extracted watermark $w_{k}$ correctly matches the expected time-dependent watermark; otherwise, the data is discarded, indicating potential tampering or an unauthorized source.
V-A1 Complexity analysis

### v2

@@ 3.3 Data Aggregation
$\sigma^{\prime}_{k}=\lambda_{k}\sum_{j=1}^{N}\delta_{j,k}=\lambda_{k}\sum_{j=1}^{N}(d_{j,k}+iw_{k})=\lambda_{k}\sigma_{k},$
with $\sigma_{k}=\sum_{j=1}^{N}(d_{j,k}+iw_{k})$. By separating the real and imaginary components of the sum, and noting that the watermark $w_{k}$ is the same for all devices in a given round, the expression for $\sigma^{\prime}_{k}$ can be rearranged as:
@@
The extraction of both the aggregated data $S_{k}$ and the watermark $w_{k}$ from $\sigma^{\prime}_{k}$ is straightforward. We can first multiply $\sigma^{\prime}_{k}$ by the inverse of $\lambda_{k}$ in the complex domain ($\lambda_{k}^{-1}$). Then, the real part of the result directly yields the aggregated original data ($S_{k}$), and the imaginary part, divided by the number of devices ($N$), gives the watermark ($w_{k}$). This extraction process can be formally expressed as:
@@
A numerical example with small numbers is provided in Section 3.5 to illustrate how the embedding and recovery methods work, both for single data and for data aggregration.
@@
3.4 Generalized Vector-based Approach for RDH
where a $2\times 2$ matrix makes the appropriate transformation.
Remark 11.
3.5 Numerical Example for Watermark Embedding, Extraction and Data Recovery
$\lambda_{1}^{-1}=\frac{3-2i}{3^{2}+2^{2}}=\frac{3}{13}-\frac{2}{13}i.$
The corresponding embedded data are the following:
Remark 12.
Remark 13.
At the receiver end, the following computation would be performed:
Thus, taking the real and the imaginary parts after the multiplication recovers both the original data (5, 8 or 17) and the watermark (4), respectively.
$S_{1}=\sum_{j=1}^{3}d_{j,1}=d_{1,1}+d_{1,2}+d_{1,3}=5+8+17=30.$
The receiver end would compute the following:
4 Homomorphic Encryption on the Complex Domain
@@
For a better understanding of the adaptation of this cryptosystem to the complex domain, a brief introduction to modular arithmetic over Gaussian integers is provided in A. A more detailed and comprehensive reference on Gaussian integers can be found in (Conrad, 2016).
4.1 ElGamal Encryption over Gaussian Integers
@@
We define the “Gaussian version” of ElGamal encryption as follows. Let $p$ be a Gaussian prime (preferably a regular prime integer of the form 4n14n-1, as discussed in A). Let $\gamma$ be a generator of the cyclic multiplicative group of Gaussian integers modulo $p$, denoted as ${\mathbb{Z}[i]}^{*}_{p}$. The ElGamal encryption scheme over Gaussian integers consists of the same three fundamental algorithms:
@@
As a key property, the order of this group, $\mathcal{N}=p^{2}-1$, is significantly larger than p1p-1 for the same numerical prime $p$ used in the regular integer ElGamal. This relationship, $\mathcal{N}=(p-1)(p+1)$, is particularly advantageous for security. If a prime $p$ is chosen such that p1p-1 possesses a large prime factor (a standard requirement for robust security in integer-based ElGamal against algorithms like Pohlig-Hellman (Pohlig and Hellman, 1978)), then this same large prime factor will inherently be a factor of $\mathcal{N}$. Therefore, the Gaussian integer ElGamal variant naturally inherits at least the same level of security (in terms of resistance to Pohlig-Hellman attacks) as its regular integer counterpart for a given $p$. Furthermore, the existence of the additional factor $(p+1)$ in $\mathcal{N}$ can potentially introduce even larger prime factors, further enhancing the security of the scheme. Thus, careful selection of the Gaussian prime $p$ to ensure that the largest prime factor of $\mathcal{N}$ is sufficiently large guides robust parameter choices.
@@ 4.1.1 Numerical example
For the parameters, the prime modulus $p=23$ and the generator $\gamma=1+2i$ are taken, and the private key is chosen as $a=7$. Now, $\mathcal{K}$ is computed as
$\mathcal{K}=\gamma^{a}\,(\operatorname{\mathrm{mod}}p)=(1+2i)^{7}\,(\operatorname{\mathrm{mod}}23)=6+2i,$
and the public key is thus formed as the triple $\mathbf{K}=(p,\gamma,\mathcal{K})=(23,1+2i,6+2i)$.
To encrypt a first message $\mu_{1}=5+4i$, an ephemeral key $b_{1}=5$ is chosen. The ciphertext components are computed as:
$\begin{split}\psi_{1}&=\gamma^{b_{1}}\,(\operatorname{\mathrm{mod}}p)=(1+2i)^{5}\,(\operatorname{\mathrm{mod}}23)=18+8i,\\ \psi_{2}&=\mu_{1}\cdot\mathcal{K}^{b_{1}}\,(\operatorname{\mathrm{mod}}p)=(5+4i)\cdot(6+2i)^{5}\,(\operatorname{\mathrm{mod}}23)=21+11i.\end{split}$
The resulting ciphertext is $(\psi_{1},\psi_{2})=(18+8i,21+11i)$. Decryption of this ciphertext using the private key $a=7$ successfully yields $\mu_{1}=(5+4i)\,(\operatorname{\mathrm{mod}}23)$.
For a second message $\mu_{2}=3+2i$, an ephemeral key $b_{2}=7$ is chosen. The ciphertext components are obtained as:
$\begin{split}\psi^{\prime}_{1}&=\gamma^{b_{2}}\,(\operatorname{\mathrm{mod}}p)=(1+2i)^{7}\,(\operatorname{\mathrm{mod}}23)=6+2i,\\ \psi^{\prime}_{2}&=\mu_{2}\cdot\mathcal{K}^{b_{2}}\,(\operatorname{\mathrm{mod}}p)=(3+2i)\cdot(6+2i)^{7}\,(\operatorname{\mathrm{mod}}23)=21+16i.\end{split}$
The resulting ciphertext is $(\psi^{\prime}_{1},\psi^{\prime}_{2})=(6+2i,21+16i)$. Decryption of this ciphertext using the private key $a=7$ successfully yields $\mu_{2}=(3+2i)\,(\operatorname{\mathrm{mod}}23)$.
@@
$\begin{split}\psi^{\prime\prime}_{1}&=\psi_{1}\cdot\psi^{\prime}_{1}=(18+8i)\cdot(6+2i)=(0+15i)\,(\operatorname{\mathrm{mod}}23),\\ \psi^{\prime\prime}_{2}&=\psi_{2}\cdot\psi^{\prime}_{2}=(21+11i)\cdot(21+16i)=(12+15i)\,(\operatorname{\mathrm{mod}}23).\end{split}$
@@
Decryption of this combined ciphertext $(\psi^{\prime\prime}_{1},\psi^{\prime\prime}_{2})=(0+15i,12+15i)$ using the private key $a=7$ yields $\mu^{\prime\prime}=(7+22i)\,(\operatorname{\mathrm{mod}}23)$. This precisely matches the direct multiplication of the original plaintexts:
$\begin{split}\mu_{1}\cdot\mu_{2}&=(5+4i)\cdot(3+2i)\,(\operatorname{\mathrm{mod}}23)\\ &=\left((15-8)+(10+12)i\right)\,(\operatorname{\mathrm{mod}}23)\\ &=(7+22i)\,(\operatorname{\mathrm{mod}}23).\end{split}$
This confirms the correct operation and homomorphic property of ElGamal encryption over Gaussian integers.
4.2 Component-wise Paillier Encryption of Gaussian Integers
While the previously discussed ElGamal scheme was extended to operate natively over Gaussian integers, the additive homomorphic property of Paillier encryption allows for a straightforward application to complex numbers (specifically, Gaussian integers) without requiring a specialized “complex version” of the cryptosystem. This is achieved by encrypting the real and imaginary parts of a Gaussian integer independently.
@@ 4.2 Component-wise Paillier Encryption of Gaussian Integers
$\mathrm{Dec}_{\mathsf{K_{priv}}}(c_{\mathrm{sum},R})$ $=(m_{1R}+m_{2R})\,(\operatorname{\mathrm{mod}}n),$
$\mathrm{Dec}_{\mathsf{K_{priv}}}(c_{\mathrm{sum},I})$ $=(m_{1I}+m_{2I})\,(\operatorname{\mathrm{mod}}n).$
@@
5 Protocols for Joint Watermarking and Encryption on the Complex Domain
@@
The first protocol, termed H[i][i]dden-EG, uses Gaussian ElGamal encryption to provide joint watermarking and encryption for individual data points. The second one, denoted as the H[i][i]dden-AggP protocol, uses Paillier encryption’s additive homomorphism to enable joint watermarking, encryption, and privacy-preserving aggregation of data from multiple sources. Each protocol is presented with its initialization and data exchange phases, and also their respective complexities.
5.1 System Model and Security Requirements
Establishing a rigorous security foundation for dispersed data in distributed environments requires a clear definition of the participating entities and the adversarial assumptions. The H[i][i]dden framework is specifically designed to address the vulnerabilities inherent in resource-constrained networks, where data must be transmitted across potentially insecure channels while maintaining perfect reversibility. In this section, we formalize the system model as a precursor to the detailed analysis of specific threat vectors and protocol-specific defenses.
5.1.1 System Entities
The H[i][i]dden framework involves the following primary entities:
  • Sensors ($\mathrm{S}_{j}$): Distributed, resource-constrained devices that generate original scalar data $d_{j,k}$. They perform the reversible embedding of the watermark $w_{k}$ and the initial encryption steps.
  • Partial Aggregators ($\mathrm{S}_{\ell}$): In the H[i][i]dden-AggP (Paillier-based) version, certain sensors or intermediate nodes act as aggregators. They use homomorphic properties to sum ciphertexts without accessing plaintext data.
  • Data Collector (DC): The central entity that initializes the system parameters, distributes challenge factors $\lambda_{k}$, and performs the final decryption to recover both the fine-grained (or aggregated) data and the watermark.
5.1.2 Security and Privacy Requirements
The following properties define the core security goals of the framework.
Definition 6 (Confidentiality).
The sensor data $d_{j,k}$ and challenge factors $\lambda_{k}$ must remain protected from unauthorized parties. This is ensured by the semantic security of the Gaussian ElGamal and Paillier encryption schemes.
Definition 7 (Integrity and Accuracy).
The system must ensure that data has not been altered during transmission or aggregation. Accuracy is guaranteed by the perfect reversibility of the complex-domain embedding, while integrity is verified through the successful extraction of the pre-agreed watermark $w_{k}$.
Definition 8 (Data Provenance).
The framework must provide proof of the origin of the data. Since the watermark $w_{k}$ is secretly agreed upon between the DC and the legitimate sensors $\mathrm{S}_{j}$, the extraction of the correct watermark serves as a cryptographic proof that the data originated from an authorized source.
Definition 9 (Privacy).
Privacy is defined as the inability of any entity (other than the DC) to learn individual sensor readings. This is maintained even during aggregation via additive homomorphic encryption.
5.1.3 Attack Model
To analyze the security of the framework, we establish the following assumptions and adversarial capabilities:
  1. a)
    Entity Assumptions: The sensors $\mathrm{S}_{j}$ are assumed to be honest-but-curious; they follow the protocol but may attempt to learn information about other sensors’ data. The Data Collertor (DC) is assumed to be a trusted entity.
  2. b)
    Adversary Capabilities: We assume a potential Attacker ($\mathcal{A}$) that can be either passive or active. $\mathcal{A}$ may reside on any communication line in the system.
  3. c)
    Threat Vectors: Within this model, the system must defend against:
    • Eavesdropping: Interception of ciphertexts to compromise confidentiality.
    • Man-in-the-Middle (MitM): Active modification of messages between entities.
    • Masquerading: Attempting to impersonate a legitimate sensor.
    • Replay Attacks: Re-sending valid messages from previous rounds to skew current data.
    • False Data Injection (FDI): Injecting malicious readings into the individual or aggregated stream.
Figure 1: Graphical representation of the H[i][i]dden-EG protocol, including the initialization (above the horizontal dashed line) and the data exchange (below the horizontal dashed line) phases.
5.2 H[i][i]dden-EG: Joint Watermarking and Encryption for Individual Data
@@
This section describes a protocol for secure and reversible information hiding between a sensor (S) and the Data Collector (DC). It applies the proposed complex domain framework and Gaussian ElGamal encryption to ensure sensor data ($d_{k}$) is transmitted confidentially while simultaneously embedding a time-dependent watermark ($w_{k}$) for verification and integrity.
Remark 14.
Although the data $d_{j,k}$ and the watermark $w_{k}$ were assumed to be simply integers (Definitions 1 and 2), the constraint of working modulo $p$ for a Gaussian prime satisfying $p\equiv 3\,(\operatorname{\mathrm{mod}}4)$ in the H[i][i]dden-EG protocol imposes further constraints on the ranges of the data, the watermark, and the challenge factor. If both positive and negative values are allowed, the Gaussian integers $(\operatorname{\mathrm{mod}}p)$ can represent any value in:
${\mathbb{Z}[i]}^{*}_{p}=\left\{x+iy\mid x,y\in\left\{-\frac{p-1}{2},-\frac{p-1}{2}+1,\dots,\frac{p-1}{2}\right\}\text{ and }(x,y)\neq(0,0)\right\}.$
Hence, the following condition is required to avoid wrap-around: $\max(|\operatorname{Re}(\lambda_{k}\delta_{k})|,|\operatorname{Im}(\lambda_{k}\delta_{k})|)<p/2.$ $\triangle$
@@
The protocol proceeds in two distinct phases: Initialization and watermarking and data exchange. The initialization phase establishes the necessary cryptographic keys and shared secrets between the DC and S:
@@
    Watermark Agreement: DC and S agree on a set of shared watermarking keys or a mechanism to generate sequential, time-dependent watermarks: $w_{1},w_{2},\ldots,w_{M}$ corresponding to times $t_{1},t_{2},\ldots,t_{M}$. Each $w_{j}$ is an integer represented with $B$ bits (e.g., taking values in the range $0,\ldots,2^{B}-1$ for unsigned watermarks, or $2^{B-1},-2^{B-1}+1,\dots,2^{B-1}-1$ for signed ones).
@@
    This step simultaneously performs both embedding and encryption, with the additional benefit of homomorphic multiplication. S then sends the resulting combined ciphertext $(\psi^{\prime\prime}_{k,1},\psi^{\prime\prime}_{k,2})$ to the DC.
@@
    DC’s Decryption and Verification: Upon receiving the combined ciphertext $(\psi^{\prime\prime}_{k,1},\psi^{\prime\prime}_{k,2})$, the DC decrypts it using its private key $a$. Due to the homomorphic property, the decrypted value $\mu_{k}=\mathrm{Dec}_{a}(\psi^{\prime\prime}_{k,1},\psi^{\prime\prime}_{k,2})$ will correspond to the product of the two original plaintexts: $\mu_{k}=\lambda_{k}\cdot\delta_{k}\,(\operatorname{\mathrm{mod}}p)$. Given the careful selection of $p$ to be sufficiently large, this modular result $\mu_{k}$ is guaranteed to be equivalent to the exact, non-modular product $\lambda_{k}\cdot\delta_{k}$ as Gaussian integers. To isolate the watermarked data, DC then computes $\delta_{k}=\mu_{k}/\lambda_{k}$ using standard complex number division. Finally, DC extracts the original sensor data $d_{k}$ from the real part of $\delta_{k}$ and the embedded watermark $w_{k}$ from its imaginary part. The data $d_{k}$ is accepted only if the extracted watermark $w_{k}$ correctly matches the expected time-dependent watermark; otherwise, the data is discarded, indicating potential tampering or an unauthorized source.
A graphical representation of the protocol is shown in Fig. 1.
5.2.1 Complexity analysis